zarkam21:

Bill forgets to put his car in park, and it starts rolling forward. When it is moving at a speed of 2 m/s, it collides with Tanya's car, which is stationary. Bill's car bounces backward at a speed of 1.2 m/s. Bill's car has a mass of 1100 kg, and Tanya's car has a mass of 800 kg. a. What is the total momentum of the system before the collision? b. What is the total momentum of the system after the collision? c. What is the velocity of Tanya's car after the collision?

NOCH:

Since Bill’s car is the only car that is moving before the collision, the total momentum is the momentum is the momentum of Bill’s car. M = 1100 * 2 = 2200 b. What is the total momentum of the system after the collision? Since momentum is conserved, the total momentum of the system after the collision is equal to the total momentum of the system before the collision. 2200 c. What is the velocity of Tanya's car after the collision? Since Bill is moving backward after the collision, his velocity is negative. Bill’s car’s final momentum = 1100 * -1.2 = -1320 To determine the velocity of Tanya's car after the collision, use the following equation. 800 * v – 1320 = 2200 800 * v = 3520 v = 3520 ÷ 800 = 4.4 m/s

zarkam21:

SO A 2200 B 2200 C 4.4 ?

zarkam21:

@Vocaloid

Vocaloid:

yeah that's right

zarkam21:

for a and b it would be kg m/s and m/s for c

Vocaloid:

good